# SALAD-Bench: A Hierarchical and Comprehensive Safety Benchmark for Large Language Models

### v3

@@ Abstract
Lijun Li1normal-$\star$, Bowen Dong1,2normal-$\star$, Ruohui Wang1normal-$\star$, Xuhao Hu1,3normal-$\star$, Wangmeng Zuo2, Dahua Lin1,4, Yu Qiao1, Jing ${}^{{\dagger}}$ ${}^{1}$ Shanghai Artificial Intelligence Laboratory ${}^{2}$ Harbin Institute of Technology ${}^{3}$ Beijing Institute of Technology ${}^{4}$ Chinese University of Hong Kong {lilijun, shaojing}@pjlab.org.cn
@@ 1 Introduction
With rapid breakthrough in LLM capabilities, new benchmarks have emerged to evaluate these models and explore their capability boundaries (Cobbe et al., 2021; Zellers et al., 2019; Clark et al., 2018; Talmor et al., 2018; Hendrycks et al., 2020). Alongside their powerful capabilities, concerns regarding the safety of LLMs are also rising. Preventing safety threats posed by generative AI systems is becoming a critical priority for both researchers (Siddiqui, 2023; Bengio, 2023; Anthropic, 2023) and policymakers (House, 2023; Kazim et al., 2023), meanwhile motivate us to explore how to comprehensively evaluate the safety capability of LLMs.
@@
SALAD-Bench (Ours) 30k H&GPT 6-16-65 Safety&Attack&Defense MD/MCQ-Judge
@@
To formulate and evaluate safety concerns, a range of benchmarks (Hartvigsen et al., 2022; Lin et al., 2023; Gehman et al., 2020; Dhamala et al., 2021; Sun et al., 2023; Wang et al., 2023b) has been developed. However, these prior benchmarks focused on safety often exhibited significant shortcomings. Firstly, most of benchmarks only focus on a narrow perspective of safety threats (e.g., only unsafe instructions or only toxic representation), failing to cover the wide spectrum of potentially harmful outputs LLMs might generate. This inadequacy partly stems from the rapid evolution of language and the emergence of new forms of harmful content, which older benchmarks failed to anticipate. Secondly, traditional harmful questions can be effectively handled with a high safety rate of about 99% by modern LLMs (Wang et al., 2023b; Sun et al., 2023). More challenging questions Chao et al. (2023); Liu et al. (2023b) are desired for comprehensive evaluation of LLM safety. Thirdly, many existing benchmarks rely on time-consuming human evaluation (Xu et al., 2023; Xie et al., 2023) or expensive GPT (Sun et al., 2023), making safety evaluation both slow and costly. Finally, these benchmarks tend to be limited in scope, being tailored either exclusively for safety evaluation (Wang et al., 2023b; Shen et al., 2023) or for testing attack and defense mechanisms (Zou et al., 2023), restricting their broader application.
@@
(1) Compact Taxonomy with Hierarchical Levels. SALAD-Bench introduces a structured hierarchy with three levels, comprising 6 domains, 16 tasks, and 65 categories, respectively. This ensures in-depth evaluation, focusing not just on overall safety but also on specific safety dimensions. As illustrated in Figure 1, a high overall safety rate does not obscure the identification of tasks and categories that may present risks. The full hierarchy of our benchmark is depicted in Figure 2.
@@
(3) Reliable and Seamless Evaluator. Leveraging instruction following capabilities, we develop two distinct evaluators for SALAD-Bench. The first, MD-Judge, is an LLM-based evaluator tailored for question-answer pairs. This model undergoes finetuning on a dataset comprising both standard and attack-enhanced pairs, labeled in alignment with our taxonomy. MD-Judge integrates relevant taxonomy details into its input and employs customized instruction tasks for precise classification. For multiple-choice question (MCQ) evaluations, we also utilize the instruction following abilities to assess the performance by regex parsing.
@@
All above contributions involve a detailed safety-focused evaluation of recent LLMs, including both black-box LLMs OpenAI (2023, 2022); Anthropic (2022); Team et al. (2023) and open-sourced models Jiang et al. (2023); Bai et al. (2023); Inan et al. (2023); Chiang et al. (2023). We analyze their vulnerabilities and assess their safety rates across different dimensions, using innovative methods to enhance the evaluation’s efficiency and scalability.
Figure 2: SALAD-Bench’s taxonomy with three levels and 65 categories focused on safety issues. Each category is represented by at least 200 questions, guaranteeing a thorough evaluation across all areas.
@@ 2 Dataset
Table 1 showcases SALAD-Bench’s advancements in LLM safety evaluations. It features 21k test samples structured into a detailed hierarchy of 6 domains, 16 tasks, and 65 categories, allowing for in-depth analysis. The benchmark is further equipped with 5k attack-enhanced and 200 defense-enhanced questions, alongside 4k multiple-choice questions, enhancing its robustness testing capabilities. Efficiency in evaluation is achieved through the use of the MD-Judge evaluator. In the following, we discuss the construction procedure.
@@ 2.1 Hierarchical Taxonomy Definition
Each domain is further subdivided into tasks and actions, resulting in 16 task-level and 65 category-level taxonomies for precise safety topic delineation. Further details on these subdivisions are provided in Appendix A.
@@ 2.2 Data Collection
Auto Labeling. To categorize questions from public datasets into SALAD-Bench’s category-level taxonomies, we employ LLMs for automated labeling through in-context learning and consensus voting. We start by crafting a template for LLM classification that outlines the task, provides few-shot learning examples, and specifies the output format, detailed in Appendix Figure 7. Using this template and a small, manually-annotated test set, we evaluate various LLMs for their labeling accuracy and select Mixtral-8x7B-Instruct (Jiang et al., 2024), Mistral-7B-Instruct (Jiang et al., 2023), and TuluV2-dpo-70B (Ivison et al., 2023) for the task. The final categorization for each question is determined by unanimous agreement among the chosen LLMs. This process ensures that every question is accurately labeled, ready for multi-dimensional analysis within the benchmark.
Following these steps, we finally obtain the base set of SALAD-Bench, containing over 200 questions per category, suitable for assessing the basic safety capabilities of LLMs. Moving forward, we will explore ways of enhancing questions to further extend the polymorphism and multifunctionality of our benchmark.
@@ 3.1 Attack Enhancement
To further explore the vulnerabilities of LLMs and examine their robustness to attacking methods, we construct the attack-enhance subset by picking harmful questions not commonly rejected by LLMs and further enhance them with attack methods. The construction steps are summarized in Figure 3.
@@
Attack Enhancement. We enhance each base question $\mathbf{q}^{\mathrm{B}}_{i}$ in $\tilde{\mathcal{Q}^{\mathrm{B}}}$ with multiple attack methods, including human designed prompts Zhu et al. (2023); Bhardwaj and Poria (2023), red-teaming LLMs Yu et al. (2023); Mehrotra et al. (2023); Liu et al. (2023a), and gradient-based methods Zou et al. (2023), and get a list of enhanced questions $\{\mathbf{q}^{\mathrm{A}}_{i,j}\}$. We list details in Appendix B.3. The final candidate set $\tilde{\mathcal{Q}^{\mathrm{A}}}$ contains ~240k questions.
@@ 3.3 Multiple-choice Questions Subset
To broaden our dataset’s scope with varied question formats, we add a multiple-choice questions (MCQ) subset, enhancing the benchmark’s complexity. Generally, we propose a heuristic self-instruct method to build the MCQ subset as follows. We first uniformly sample $K$ harmful questions in each category. For each question $\mathbf{q}$, we generate multiple safe responses and unsafe responses as candidates. The safe and unsafe responses are generated using GPT-4 (OpenAI, 2023) and fine-tuned GPT in Section 2.2 respectively. Then, after selection parsing and human re-checking, we randomly select totally three safe and unsafe selections and obtain final choices $\mathcal{S}$. Finally, for each $\mathbf{q}$ with $\mathcal{S}$, we construct two questions (i.e., choosing safe or unsafe choices) by prompting method. All the templates are discussed in Appendix H. Finally, we obtain the target MCQ subset with ~4k questions and covers all categories listed in Section 2.1.
@@ 4.1 Evaluation for QA-Pairs
Previous LLM safety benchmarks (Zou et al., 2023; Deng et al., 2023) usually utilize 1) keyword matching (Zou et al., 2023), 2) moderation classifiers (Hanu and Unitary team, 2020) and 3) GPT (Deng et al., 2023) to conduct safety evaluation. However, keyword matching-based evaluator struggles with various outputs from LLMs; moderation classifiers only focus on a narrow spectrum of safety threats; GPT incurs much extra cost to call APIs, and upgrade of the backend model affects reproducing evaluation results.
Figure 4: An example of our training sample with domain-level taxonomy as safety category.
Inspired by LlamaGuard (Inan et al., 2023), we use public and self-generated data to fine-tune an LLM-based classifier from Mistral-7B (Jiang et al., 2023). Distinctly, our task involves evaluating not only plain question-answer pairs but also attack-enhanced question-answer pairs. Our evaluator is named MD-Judge.
To make our MD-Judge capable of both plain and attack-enhanced questions, we collect plain QA pairs from previous works (Ji et al., 2023; Zheng et al., 2023a; Lin et al., 2023) and construct both safe and unsafe answers to enhanced questions. The safety labels of attack-enhanced QA pairs are labeled by GPT-4. Finally, we utilize the Auto Labeling toolkit illustrated in Section 2.2 to annotate all training samples within the taxonomies of SALAD-Bench. During fine-tuning, we propose a safety evaluation template to reformat question-answer pairs for MD-Judge predictions, as shown in Figure 4. Besides, the template’s structure are detailed described in Appendix D. This structured data then undergoes fine-tuning to enhance MD-Judge’s capabilities.
@@ 4.2 Evaluation for Multiple-choice Questions
Previous works use keyword-based evaluators (Xu et al., 2023) and GPT-based evaluators Deng et al. (2023) to handle open-ended responses for multiple-choice questions (MCQ). To sustain high evaluation accuracy meanwhile reduce the cost from inference, we introduce MCQ-Judge, which leverages in-context learning with regex parsing to efficiently fetch the answers. Specifically, we first define <ans> token as well as <eoa> token to strictly wrap the output selections, and obtain the basic instruction of evaluation prompt. Then, to leverage the instruction following ability of LLMs and obtain the formatted responses, we manually construct few-shot exemplars as prompts to conduct in-context learning. Hence we obtain the final prompt for MCQ-Judge, and the total prompts are listed in Appendix H.
@@ 5.1 Experiment settings
Attack methods. We evaluate TAP Mehrotra et al. (2023), AutoDAN Liu et al. (2023a), GPTFuzz Yu et al. (2023), GCG Zou et al. (2023), CoU Bhardwaj and Poria (2023) and human designed jailbreaks. For TAP, we employ vicuna-13B-v1.5, GPT-3.5 and GPT-4 as the evaluator. For AutoDAN, we use GPT-3.5 as mutator or do not use mutator. For GCG we follow Robey et al. (2023) to use 20 beforehand searched suffixes. All methods use Llama-2-7B-Chat as the target model. More details are listed in Appendix G.
@@
Table 3: Comparison of F1 scores between our model and other leading methods. Best results are bolded and second best are underlined [^2] Same in the following tables.. Base and Enhance indicate our SALAD-Base-Test and SALAD-Enhance-Test, TC means ToxicChat, and Beaver means Beavertails.
5.2 Evaluator Comparison
We fine-tune MD-Judge from Mistral-7B (Jiang et al., 2023) with sequence length of 4096 via LoRA (Hu et al., 2021) with Rank=$64$. The model underwent training on 8xA100 GPUs, with a per-GPU batch size of 16, over a total of 2 training epochs. Furthermore, we also fine-tune on different base models to compare the performances, which is shown in Appendix F.
Test Dataset. We test MD-Judge on several datasets, including self-generated&annotated and publicly available ones. We create SALAD-Base-Test and SALAD-Enhance-Test from SALAD-Bench to represent base and attack-enhanced test sets to assess different evaluators. We also use public test sets, i.e., ToxicChat (Lin et al., 2023), Beavertails (Ji et al., 2023), and a 2k subset of SafeRLHF (Dai et al., 2023) test set for evaluation.
Implementation Details. We compare MD-Judge with three methods, i.e., keywords evaluator, gpt-based evaluators (gpt-3.5-turbo-1106 and gpt-4-1106-preview), and LlamaGuard Inan et al. (2023).
Evaluator Results.
Evaluation results of F1 scores are detailed in Table 3. The comprehensive outcomes reveal that MD-Judge markedly surpasses its counterparts on both our proprietary test set and the publicly available safety test set, underscoring the effectiveness of MD-Judge’s fine-tuning for enhanced general safety guard capabilities. For more in-depth results and analysis, please refer to Appendix E.

### v4

@@ Abstract
[…]
@@ 1 Introduction
With rapid breakthrough in LLM capabilities, new benchmarks have emerged to evaluate these models and explore their capability boundaries (Cobbe et al., 2021; Zellers et al., 2019; Clark et al., 2018; Talmor et al., 2018; Hendrycks et al., 2020). Alongside their powerful capabilities, concerns regarding the safety of LLMs are also rising. Preventing safety threats posed by generative AI systems is becoming a critical priority for both researchers (Siddiqui, 2023; Bengio, 2023; Anthropic, 2023) and policymakers (House, 2023; Kazim et al., 2023), meanwhile motivates us to explore how to comprehensively evaluate the safety capability of LLMs.
@@
SALAD-Bench (Ours) 30k H&GPT 6-16-66 Safety&Attack&Defense MD/MCQ-Judge
@@
To formulate and evaluate safety concerns, a range of benchmarks (Hartvigsen et al., 2022; Lin et al., 2023; Gehman et al., 2020; Dhamala et al., 2021; Sun et al., 2023; Wang et al., 2023b) have been developed. However, these prior benchmarks focused on safety often exhibited significant shortcomings. Firstly, most of benchmarks only focus on a narrow perspective of safety threats (e.g., only unsafe instructions or only toxic representation), failing to cover the wide spectrum of potentially harmful outputs LLMs might generate. This inadequacy partly stems from the rapid evolution of language and the emergence of new forms of harmful content, which older benchmarks failed to anticipate. Secondly, traditional harmful questions can be effectively handled with a high safety rate of about 99% by modern LLMs (Wang et al., 2023b; Sun et al., 2023). More challenging questions Chao et al. (2023); Liu et al. (2023b) are desired for comprehensive evaluation of LLM safety. Thirdly, many existing benchmarks rely on time-consuming human evaluation (Xu et al., 2023; Xie et al., 2023) or expensive GPT (Sun et al., 2023), making safety evaluation both slow and costly. Finally, these benchmarks tend to be limited in scope, being tailored either exclusively for safety evaluation (Wang et al., 2023b; Shen et al., 2023) or for testing attack and defense mechanisms (Zou et al., 2023), restricting their broader application.
@@
(1) Compact Taxonomy with Hierarchical Levels. SALAD-Bench introduces a structured hierarchy with three levels, comprising 6 domains, 16 tasks, and 66 categories, respectively. This ensures in-depth evaluation, focusing not just on overall safety but also on specific safety dimensions. As illustrated in Figure 1, a high overall safety rate does not obscure the identification of tasks and categories that may present risks. The full hierarchy of our benchmark is depicted in Figure 2.
@@
(3) Reliable and Seamless Evaluator. Leveraging instruction following capabilities, we develop two distinct evaluators for SALAD-Bench. The first, MD-Judge, short for Multi-Dimension Judge, an LLM-based evaluator tailored for question-answer pairs. This model undergoes finetuning on a dataset comprising both standard and attack-enhanced pairs, labeled in alignment with our taxonomy. MD-Judge integrates relevant taxonomy details into its input and employs customized instruction tasks for precise classification. For multiple-choice question (MCQ) evaluations, we also utilize the instruction following abilities to assess the performance by regex parsing, which we called as MCQ-Judge.
@@
All above contributions involve a detailed safety-focused evaluation of recent LLMs, including both black-box LLMs OpenAI (2023a, 2022); Anthropic (2022); Team et al. (2023) and open-sourced models Jiang et al. (2023); Bai et al. (2023); Touvron et al. (2023); Chiang et al. (2023). We analyze their vulnerabilities and assess their safety rates across different dimensions, using innovative methods to enhance the evaluation’s efficiency and scalability.
Figure 2: SALAD-Bench’s taxonomy with three levels and 66 categories focused on safety issues. Each category is represented by at least 200 questions, guaranteeing a thorough evaluation across all areas.
@@ 2 Dataset
Table 1 showcases SALAD-Bench’s advancements in LLM safety evaluations. It features 21k test samples structured into a detailed hierarchy of 6 domains, 16 tasks, and 66 categories, allowing for in-depth analysis. The benchmark is further equipped with 5k attack-enhanced and 200 defense-enhanced questions, alongside 4k multiple-choice questions, enhancing its robustness testing capabilities. Efficiency in evaluation is achieved through the use of the MD-Judge evaluator. In the following, we discuss the construction procedure.
@@ 2.1 Hierarchical Taxonomy Definition
Each domain is further subdivided into tasks and actions, resulting in 16 task-level and 66 category-level taxonomies for precise safety topic delineation. Further details on these subdivisions are provided in Appendix B.
@@ 2.2 Data Collection
Auto Labeling. To categorize questions from public datasets into SALAD-Bench’s category-level taxonomies, we employ LLMs for automated labeling through in-context learning and consensus voting. We start by crafting a template for LLM classification that outlines the task, provides few-shot learning examples, and specifies the output format, detailed in Appendix Figure 7. Using this template and a small, manually-annotated test set, we evaluate various LLMs for their labeling accuracy and select Mixtral-8x7B-Instruct (Jiang et al., 2024), Mistral-7B-Instruct (Jiang et al., 2023), and TuluV2-dpo-70B (Ivison et al., 2023) for the task. The final categorization for each question is determined by unanimous agreement among the chosen LLMs. This process ensures that every question is accurately labeled, ready for multi-dimensional analysis within the benchmark. To ensure the labeling results are convincing, we also conduct human verification on randomly sampled examples. During human verification, three human annotators are involved by labeling and cross-checking to obtain convincing human labels as ground-truth. The consistency rate between auto labeling and human labels achieves 94.3%.
Following these steps, we finally obtain the base set of SALAD-Bench, containing over 200 questions per category, suitable for assessing the basic safety capabilities of LLMs. To ensure the quality of our benchmark dataset, we conduct a human-verification experiment, which is detailed described in Appendix E. Moving forward, we will explore ways of enhancing questions to further extend the polymorphism and multifunctionality of our benchmark.
@@ 3.1 Attack Enhancement
To further explore the vulnerabilities of LLMs and examine their robustness to attacking methods, we construct the attack-enhance subset by picking harmful questions not commonly rejected by LLMs and further enhancing them with attack methods. The construction steps are summarized in Figure 3.
@@
Attack Enhancement. We enhance each base question $\mathbf{q}^{\mathrm{B}}_{i}$ in $\tilde{\mathcal{Q}^{\mathrm{B}}}$ with multiple attack methods, including human designed prompts Zhu et al. (2023); Bhardwaj and Poria (2023), red-teaming LLMs Yu et al. (2023); Mehrotra et al. (2023); Liu et al. (2023a), and gradient-based methods Zou et al. (2023), and get a list of enhanced questions $\{\mathbf{q}^{\mathrm{A}}_{i,j}\}$. We list details in Appendix F. The final candidate set $\tilde{\mathcal{Q}^{\mathrm{A}}}$ contains ~240k questions.
@@ 3.3 Multiple-choice Questions Subset
To broaden our dataset’s scope with varied question formats, we add a multiple-choice questions (MCQ) subset, enhancing the benchmark’s complexity. Generally, we propose a heuristic self-instruct method to build the MCQ subset as follows. We first uniformly sample $K$ harmful questions in each category. For each question $\mathbf{q}$, we generate multiple safe responses and unsafe responses as candidates. The safe and unsafe responses are generated using GPT-4 (OpenAI, 2023a) and fine-tuned GPT in Section 2.2 respectively. Then, after selection parsing and human re-checking, we randomly select totally three safe and unsafe selections and obtain final choices $\mathcal{S}$. Finally, for each $\mathbf{q}$ with $\mathcal{S}$, we construct two questions (i.e., choosing safe or unsafe choices) by prompting methods. All the templates are discussed in Appendix L. Finally, we obtain the target MCQ subset with ~4k questions and cover all categories listed in Section 2.1.
@@ 4 Evaluator
In order to boosting the development of LLM safety evaluation, we propose a reliable and seamless evaluator, designed specifically for our benchmark. The evaluator is composed of the MD-Judge and MCQ-Judge, which were constructed for question-answer pairs and multiple choice questions, respectively.
@@ 4.1 Evaluation for QA-Pairs
Considering the existing challenges mentioned in Appendix A.3, we use public and self-generated data to fine-tune an LLM-based safety judge model from Mistral-7B (Jiang et al., 2023), inspired by LlamaGuard (Inan et al., 2023). Distinctly, our task involves evaluating not only plain question-answer pairs but also attack-enhanced question-answer pairs. The training set for our MD-Judge mainly composed two sources: open-sourced safety data and specially generated data which are described in Appendix G.
Figure 4: An example of our evaluator’s template with domain-level areas as safety taxonomy.
To make our MD-Judge capable of both plain and attack-enhanced questions, we collect plain QA pairs from training set of previous works (Ji et al., 2023; Zheng et al., 2023a; Lin et al., 2023) and construct both safe and unsafe answers to enhanced questions. The safety labels of attack-enhanced QA pairs are labeled by GPT-4. Finally, we utilize the Auto Labeling toolkit illustrated in Section 2.2 to annotate all training samples within the taxonomies of SALAD-Bench.
During fine-tuning, we propose a safety evaluation template to reformat question-answer pairs for MD-Judge predictions, as shown in Figure 4. Besides, the template’s structure are detailed described in Appendix H. This structured data then undergoes fine-tuning to enhance MD-Judge’s capabilities.
@@ 4.2 Evaluation for Multiple-choice Questions
In contrast to the Question-Answering subset where ground-truth answers are provided, the multiple-choice subset includes predefined correct answers. However, due to variations in instruction-following capabilities among different LLMs, evaluation performance requires matching open-ended responses to predefined answer choices. To sustain high evaluation accuracy meanwhile reduce the cost from inference, we introduce MCQ-Judge, which leverages in-context learning with regex parsing to efficiently fetch the answers. Specifically, we first define <ans> token as well as <eoa> token to strictly wrap the output selections, and obtain the basic instruction of evaluation prompt. Then, to leverage the instruction following ability of LLMs and obtain the formatted responses, we manually construct few-shot exemplars as prompts to conduct in-context learning. Hence we obtain the final prompt for MCQ-Judge, and the total prompts are listed in Appendix L.
@@ 5 Experiments
Leveraging our high-quality datasets, we conducted large-scale experiments to assess both the reliability of our evaluators and the safety of various Large Language Models (LLMs) and compare the effectiveness of different attack and defense methods.
@@ 5.1 Experiment settings
5.1.1 Settings for Evaluator
Test Dataset. We test MD-Judge on several datasets, including self-generated and publicly available ones. We create SALAD-Base-Test and SALAD-Enhance-Test from SALAD-Bench to represent base and attack-enhanced test sets to assess different evaluators. We also use public test sets, i.e., ToxicChat (Lin et al., 2023), Beavertails (Ji et al., 2023), and a 2k subset of SafeRLHF (Dai et al., 2023) test set for evaluation.
Implementation Details. We fine-tune MD-Judge from Mistral-7B (Jiang et al., 2023) with sequence length of 4096 via LoRA (Hu et al., 2021) with Rank=$64$. The model underwent training on 8xA100 GPUs, with a per-GPU batch size of 16, over a total of 2 training epochs. Furthermore, we also fine-tune on different base models to compare the performances, which is shown in Appendix J.
5.1.2 Settings for SALAD-Bench
@@
Attack methods. We evaluate TAP Mehrotra et al. (2023), AutoDAN Liu et al. (2023a), GPTFuzz Yu et al. (2023), GCG Zou et al. (2023), CoU Bhardwaj and Poria (2023) and human designed jailbreaks. For TAP, we employ vicuna-13B-v1.5, GPT-3.5 and GPT-4 as the evaluator. For AutoDAN, we use GPT-3.5 as mutator or do not use mutator. For GCG we follow Robey et al. (2023) to use 20 beforehand searched suffixes. All methods use Llama-2-7B-Chat as the target model. More details are listed in Appendix M.
@@
Table 3: Comparison of F1 scores between MD-Judge and other leading methods. Best results are bolded and second best are underlined [^2] Same in the following tables.. Base and Enhance indicate our SALAD-Base-Test and SALAD-Enhance-Test, TC means ToxicChat, and Beaver means Beavertails.
  Methods Harmbench Lifetox
  GPT-3.5 61.13% 73.39%
GPT-4 84.46% 77.43%
LlamaGuard 64.56% 58.62%
MD-Judge 83.72% 79.27%
Table 4: Comparison of accuracy between MD-Judge and other leading methods on out-of-distributed Harmbench and Lifetox benchmarks. Best results are bolded and second best are underlined. MD-Judge achieves comparable performance against state-of-the-art methods on out-of-distributed safety benchmarks.
[…]
5.2 Evaluation of MD-Judge
Comparison with other evaluators.
We compare MD-Judge with three methods, i.e., keywords evaluator, gpt-based evaluators (gpt-3.5-turbo-1106 and gpt-4-1106-preview), and LlamaGuard Inan et al. (2023). Evaluation results of F1 scores are detailed in Table 3. The comprehensive outcomes reveal that MD-Judge markedly surpasses its counterparts on both our proprietary test set and the publicly available safety test set, underscoring the effectiveness of MD-Judge’s fine-tuning for enhanced general safety guard capabilities. For more in-depth results and analysis, kindly refer to Appendix I for details. Moreover, we also conduct evaluation on the out-of-distribution Harmbench (Mazeika et al., 2024) and Lifetox (Kim et al., 2023) benchmark to further investigate the robustness of MD-Judge and the counterparts. Table 4 shows the evaluation results on both datasets. MD-Judge achieves comparable performance against the state-of-the-art GPT-4 evaluator, and largely surpasses LlamaGuard Inan et al. (2023) by at least 20%. Above promising results prove that our MD-Judge has reliable evaluation capability in safety aspects.
5.3 Evaluation of MCQ-Judge
In terms of the MCQ-Judge comparison experiments, we compare MCQ-Judge with different previously proposed evaluators (i.e., keyword-based and GPT-based evaluators), and introduce human evaluator as reference. Experimental results show that MCQ-Judge performs much closer to human evaluation results than any other counterpart with only 0.43s latency. Such results demonstrate the efficiency and effectiveness of MCQ-Judge. The detailed experiments settings and results are shown in Appendix K.